Consider the following: in all respects, company a and company b are identical except company a’s costs are mostly variable. Company b’s costs are mostly fixed. When sales increase, which company will tend to realize the greatest increase in profits? Explain. Should be around 200-250 words.
Answer #1

The scenario presented in the question there are two identical companies A and B and the difference between them are the company A costs are mostly variable while the company B costs are mostly fixed. When there is an increase in sales we tend to see the company with Higher fixed costs will reap more benefits and higher increase in profits than the one with fixed costs since the sales increase will not affect fixed costs as these are constant and one with more variable costs will have lower contribution margin and in turn lower net profits. The one with higher fixed costs means it is more leveraged means it is more sensitive to change in net income with change in dollar sales

Contribution margin= Sales-variable costs

Net profit=Contribution margin-Fixed costs

Let us take an example below:

Company A:

The no of units sold =1000, selling price=10 variable costs=5 and fixed costs=1000

Contribution margin=1000*(10-5)= 5000

Net income=5000-1000=4000

Company B:

The no of units sold =1000, selling price=10 variable costs=2 and fixed costs=4000

Contribution margin=1000*(10-2)=8000

Net income=8000-4000=4000

Now assume increase in sales units=2000

Company A:

Contribution margin=2000*(10-5)= 10000

Net income=10000-1000=9000

Company B:

Contribution margin=2000*(10-2)= 16000

Net income=16000-4000=12000

We can observe here when there is increase in sales the company B is having higher net income than company A